OpenStudy (help!!!!):

Mike weights x-3, his brother Tim weights 8 more pounds that mike. While his (Sarah)sister weights 12 pounds than tim. Find their total weight.

OpenStudy (help!!!!):

a.4x-5 b.5x+3 c.x+12 d.3x-5 e.2x+5

OpenStudy (anonymous):

Let M= x-3 Let T=x+5 Let S=x+17 add them together and get your answer